OpenStudy (anonymous):

Solve for x: 3x-5=2x+6 A. 1 B. -1 C. 11 D. -11

OpenStudy (anonymous):

1) Pass the left to the right, you get: 3x-5-2x-6 =0. 2) Solve the like terms, (3x-2x = x) (-5-6 =-11) 3) Now you have x-11=0. 4) Add 11 on both sides to solve for x (x-11+11=0+11). 5) X=11

OpenStudy (anonymous):

*Sorry, first step is right to left, same thing though.